### v5

@@ Results
This section aims to enhance AFLoc’s clinical applicability by presenting results obtained through leveraging limited labeled data for model fine-tuning on both segmentation and localization tasks.
As expected, AFLoc demonstrates enhanced segmentation performance with increased fine-tuning using labeled data on the SIIM dataset (Extended Data Table 9). The Dice score is increased from 0.772 with 1% of the data for fine-tuning to 0.809 with 100% of the data. Furthermore, AFLoc consistently achieves higher Dice scores than all comparative methods across varying portions of labeled fine-tuning data. These results indicate that AFLoc can effectively utilize limited annotations to achieve enhanced segmentation performance.
The results for pathology localization on the NIH ChestXray14 dataset are depicted in Extended Data Table 10. Consistent performance enhancement is achieved by AFLoc over the two comparative methods across different IoU thresholds. At an IoU of 0.3, AFLoc attains a mean localization accuracy of 0.62, surpassing the performance of the two comparative methods (Li et al. (0.49) and Wang et al. (0.22)). Notably, the model excels even under stricter thresholds like IoU = 0.5 and IoU = 0.7, especially for more challenging pathologies such as atelectasis and infiltration.
Extended Data Fig. 7 provides visual illustrations of AFLoc’s localization results on pneumothorax chest X-ray images and retinal hemorrhages. The model’s predictions align well with the ground truth, even for small and subtle lesions. These results demonstrate that AFLoc is capable of leveraging limited labeled data to improve its localization performance.
@@
In this section, we explore the impact of incorporating text features at different levels of granularity – word-level, sentence-level, and report-level – on the performance of AFLoc. The results are listed in Extended Data Table 11. When using only one level of text features, the performance of AFLoc deteriorates a lot, especially for word-level and report-level features. One possible explanation could be that the word-level features lack sufficient contextual information to capture the characteristics of the pathology, whereas the report-level features may fail to capture critical diagnostic details due to their overly macroscopic perspective. On the other hand, AFLoc using sentence-level features achieves relatively higher performance, suggesting that sentence-level features can provide richer contextual information, which is important for pathology localization in medical imaging. When these different levels of text features are combined in utilization, clear performance enhancement is observed, and the best performance is achieved when all three levels are employed, fully supporting our hypothesis that the combination of multi-level information is crucial for improving the accuracy of pathology localization in medical images.
@@
Text encoders play a crucial role in deeply understanding medical concepts and achieving accurate semantic alignment. To investigate the impact of text encoding on pathology localization performance, we implemented the text encoder from the latest large language and vision assistant for biomedicine (LLaVA-Med [46]). LLaVA-Med has shown improvements in open biomedical question-answering tasks. However, as demonstrated in Extended Data Table 12, its performance falls short compared to BioClinicalBERT [47]. This discrepancy can be attributed to the superior performance of language models trained on domain-specific datasets (e.g., biomedical and clinical data) tailored for their respective tasks [48]. LLaVA-Med was trained on the PMC-15M dataset, which is extracted from scientific publications in PubMed Central, whereas BioClinicalBERT was trained on electronic health records of intensive care unit patients, closely aligning with the clinical reports used in this study. This domain-specific alignment likely provides BioClinicalBERT an advantage in comprehending intricate medical terminology and improving its localization performance.
@@ Discussion
Pathology localization and diagnosis are essential components of clinical applications in medical imaging. However, existing methods encounter various challenges, notably stemming from the scarcity of annotated data related to generalization [19, 18]. In this study, we introduce AFLoc, an annotation-free deep learning model designed to enhance pathology localization by utilizing both clinical reports and imaging data in the pre-training phase.
@@
AFLoc employs a multi-level semantic alignment strategy to enhance the medical feature representation. Given that pathologies often occupy only small regions within images, accurate localization can be challenging for global models. While some methods align global and local text-image features [30], they may miss critical semantic context by focusing excessively on individual words. In contrast, AFLoc adopts a more comprehensive approach by incorporating features at different levels of granularity (word-level, sentence-level, and report-level) during pre-training. This allows AFLoc to capture the necessary details for precise pathology localization.
@@
We extensively evaluated AFLoc’s localization and diagnostic capabilities in an annotation-free setting across three major clinical imaging modalities: chest X-rays, retinal fundus, and histopathology. In the chest X-ray modality, AFLoc demonstrated improved localization performance compared to existing methods, such as MedKLIP and BioViL, across multiple evaluation metrics, including Intersection over Union (IoU), Dice similarity coefficient, and Contrast-to-Noise Ratio (CNR). AFLoc effectively localized a variety of thoracic pathologies, including pneumonia, atelectasis, and pleural effusion, on datasets such as RSNA Pneumonia, MS-CXR, and CheXlocalize. Notably, AFLoc showed good performance in localizing subtle and low-contrast pathologies, such as airspace opacity and atelectasis, which are often difficult to detect but crucial for early diagnosis and treatment planning. Furthermore, AFLoc surpassed the human benchmark on the CheXlocalize dataset across five key tasks, implying its potential to compete with conventional methods that depend on human-reviewed data. Additionally, AFLoc performed well on the COVID Rural dataset, which provides X-rays of patients affected with COVID-19 pneumonia, highlighting its potential for generalizing to new and emerging diseases. This capability is particularly important for timely and accurate disease detection, especially during global health crises like the COVID-19 pandemic.
@@
The utilization of AFLoc with retinal fundus images underscores its versatility. In retinal pathology, AFLoc performed well at localizing and identifying features such as choroidal neovascularization and drusen, surpassing other methods like GLoRIA. Given the subtlety and diffuseness of retinal diseases, precise localization is particularly challenging. AFLoc’s adept performance without human-annotated data highlights its potential to improve ophthalmic diagnostics, especially in settings where large-scale data annotation is not feasible.
@@
Finally, we illustrated that AFLoc’s performance could be further improved with a small amount of annotated data, demonstrating its scalability and adaptability for clinical applications. While annotation-free localization is a valuable property, fine-tuning with labeled data can elevate AFLoc’s performance. This flexibility renders AFLoc fitting for diverse clinical environments, spanning well-resourced hospitals with access to large labeled datasets to smaller clinics in resource-limited settings. AFLoc’s adeptness in achieving commendable results with minimal annotated data makes it a potentially practical solution for healthcare systems globally, particularly in regions where access to expert annotators is limited.
@@
Despite these encouraging results, AFLoc does exhibit certain limitations. There is still room for improvement, particularly in refining more precise pathological segmentation. Implementing advanced image feature extraction techniques, like aligning local features of suggested regions more effectively with text features from multi-scale anchor boxes, may offer avenues to boost its accuracy. Additionally, AFLoc’s performance could benefit from access to larger, more diverse datasets. We view AFLoc as a potential benchmark for annotation-free methodologies in the medical field, with the possibility of playing a critical role in supporting healthcare systems during forthcoming disease outbreaks.
@@ Methods
Text encoding In clinical practice, medical texts may originate from various types of content, including clinical observations, diagnostic descriptions, and related information. These texts are typically composed of multiple sentences or phrases. In GLoRIA [30], a block tokenization technique was used, constructing complete words by aggregating multiple subwords. On the other hand, BioViL [28] employs a custom dictionary developed from multiple datasets, aiming to reduce the frequency of words being split into subwords. A radiological finding in an image might require information of location, size, and severity for a comprehensive description. Treating words as independent entities may overlook the complete semantics, leading to incorrect semantic alignment between text and image. In our proposed AFLoc, text features are extracted at three levels: word-level, sentence-level, and report-level, aiming to represent report content more comprehensively and precisely through multi-granularity semantics. Additionally, this three levels structure is naturally present in the text, requiring no complex designs to accomplish our text encoding.
Specifically, we employ BioClinicalBERT [47] as our text encoder to extract text features. When processing a medical report containing words and sentences, each word is tokenized into tokens. The tokenizer inputs the tokenized report into the text encoder , where H is the maximum token length. The text encoder outputs , L is the number of layers, and indicates the feature dimension. We take the average feature of the last 4 layers as the subword-level feature . Then, we aggregate features at the word, sentence, and report levels. Particularly, for a word, all its subword-level features are summarized to obtain the word-level features . On the other hand, all corresponding subword-level features are averaged to obtain sentence-level features and one report-level feature .
[…]
@@
Annotation-free Pathology Localization and Diagnosis Once the model is trained, it can be applied to the pathology localization and clinical diagnosis tasks without requiring additional annotations. Instead, we transform the task into a text-image matching problem using text descriptions, offering greater flexibility as the text descriptions can be easily generated to meet various clinical requirements. By default, we use the automatically generated descriptions based on the pathology’s name and a template for testing. That is, for example, ‘Findings suggesting {lesion}’ for a positive description and ‘No findings suggesting {lesion}’ for a negative description. We refer to such description as ”Simple description”. In addition, AFLoc also supports the ”Precise description” of the pathology, such as ‘Large right-sided pneumothorax, air space opacity in a right infrahilar location’. This ”precise description” comes from clinicians who, based on their observation of the images, write corresponding descriptions to validate their hypotheses when the diagnosis is uncertain. With the text descriptions, the text features can be obtained with the pre-trained text encoder.
@@
Experimental details We pre-train AFLoc on MIMIC-CXR, Retinal Fundus, and Quilt-1M dataset for chest X-ray, retinal fundus, and histopathology respectively. All images were uniformly resized to pixels and we followed the same image preprocessing strategy as [30].
A random shuffling and a sentence sampling strategy were applied to enhance the text data. Notably, the reports for the chest X-ray pretrain include multiple sections with different contents. We use the texts within the ’Findings’ and ’Impressions’ sections to train our model. Furthermore, we applied an additional sentence augmentation strategy to enhance the diversity of medical reports for retinal fundus data. Specifically, we generate five semantically equivalent but differently expressed sentences for each sentence by ChatGPT. These equivalent sentences will be randomly replaced during pre-training. For the histopathology dataset, we directly use the corresponding caption provided by the dataset for pre-training.
During pretraining, we used the Adam optimizer with an initial learning rate of 0.00002. A linear learning rate decay mechanism is employed by a decay factor of 0.9. The other hyperparameters were optimized with a corresponding validation dataset.
Evaluation metrics For the localization task, the Intersection over Union (IoU) and Dice coefficients were used to evaluate the thresholded results. We report the average results at multiple thresholds following the settings from BioViL. For the CheXlocalize dataset, we followed the methodology from [14] that determined an optimal threshold from the validation set and applied it to the test set. Additionally, we used the CNR [28] for measuring the difference in heatmap scores between inside and outside the bounding box area, eliminating the need for a hard threshold. This assessment of local similarity is significant, as certain clinical downstream applications might benefit more from heatmap visualizations than discrete segmentation masks. To directly assess localization performance without the influence of classification errors, we reported localization performance on the positive slices of the dataset (those with ground truth segmentations or bounding boxes annotated by radiologists). This approach ensured that the evaluation focused on the accuracy of localization, without being confounded by classification errors [14]. The averages of these metrics are reported on the test set, based on over 1,000 bootstrap repetitions, accompanied by the 95% confidence interval (CI), which was determined using the 2.5th and 97.5th percentiles of the empirical distribution.
@@
In our chest X-ray experiments, we utilized the MIMIC-CXR dataset for training. We then evaluated the model across six external datasets: RSNA, MS-CXR, CheXlocalize, COVID Rural, SIIM-ACR Pneumothorax, and ChestX-Det10. Additionally, to validate the model’s adaptability to different modalities, we collected retinal fundus data. For the retinal fundus experiments, the model was retrained using image-report pairs from this dataset.
@@
We evaluated our model across three different imaging modalities: chest X-ray, retinal fundus, and histopathology. For the chest X-ray experiments, we pre-trained our model on the MIMIC-CXR dataset, then evaluated it across seven widely used external datasets: RSNA, MS-CXR, CheXlocalize, COVID Rural, SIIM-ACR Pneumothorax, NIH ChestXray14, and ChestX-Det10. For the retinal fundus experiments, we pre-trained and evaluated the model using a real-world in-house dataset. For the histopathology experiments, we pre-trained our model on the Quilt-1M dataset and then evaluated it on three external datasets: SICAPv2, WSSS4LUAD, and DHMC LUAD. The details of each dataset are provided below.
@@
MIMIC-CXR [50]. The MIMIC-CXR dataset is an extensive, publicly available collection of chest X-ray images, paired with free-text radiological reports. It contains 377,110 images that are linked to 227,835 radiological studies. Each study within the database may consist of one or more images, providing a diverse range of scanning perspectives for analysis and research purposes.
@@
RSNA Pneumonia [35]. The RSNA Pneumonia dataset presents a challenge in identifying visual signals of pneumonia from medical images. Pneumonia typically presents as one or multiple regions of increased opacity in chest X-rays [51]. However, diagnosing pneumonia based on chest X-rays can be complicated due to the presence of many other lung conditions, such as pulmonary edema, bleeding, atelectasis, lung cancer, or surgical changes. Additionally, pleural effusion outside the lungs can also appear as increased opacity in CXR. In this context, developing automated methods for pneumonia localization can be highly beneficial in the clinical environment. Images in this dataset are accompanied by bounding box labels that indicate the areas affected by pneumonia. Over 1000 cases in this dataset are positive, with each image potentially containing one or more boundary boxes highlighting the pneumonia regions.
@@
MS-CXR [28]. MS-CXR is a curated database that features 1,153 image-sentence pairs of bounding boxes along with their corresponding phrases. This collection encompasses 8 different types of cardiopulmonary radiological findings. The image-sentence pairs are distributed equally among the 8 findings. In addition to brief descriptions, this dataset provides authentic depictions of radiological findings as described in radiology reports and oral records. All the data have been annotated and validated by professional radiologists, ensuring a gold standard for phrase grounding assessment. MS-CXR is a subset of the MIMIC-CXR dataset. All data in MS-CXR have been removed from the MIMIC-CXR training set to ensure it was not used in the training process.
@@
CheXlocalize [14]. The CheXlocalize dataset is a comprehensively curated chest X-ray dataset. It features multi-label, pixel-level segmentations meticulously drawn by two board-certified radiologists, ensuring a highly precise ground truth annotation for each image. This dataset includes 234 images, encompassing 643 segmented pathologies. It covers a broad range of cardiopulmonary conditions, including airspace opacity, atelectasis, cardiomegaly, consolidation, edema, enlarged cardiomediastinum, lung lesion, pleural effusion, pneumothorax, and support devices. Moreover, CheXlocalize also provides human benchmark segmentations, drawn by a separate group of 3 board-certified radiologists, which can be used to study human benchmark localization performance.
@@
COVID Rural [36, 37]. The COVID Rural project includes 221 chest X-ray images from 105 patients who tested positive for COVID-19. In these images, opaque areas associated with COVID-19 infection are delineated by polygons that connect specific points. The manual annotations of these images were generated by primary physicians and reviewed by experienced radiologists.
@@
SIIM-ACR Pneumothorax (SIIM) [52]. The SIIM dataset is a widely adopted dataset for the evaluation of chest X-ray models. This dataset is designed to promote the development of models for identifying pneumothorax. In severe cases, pneumothorax can cause obstructive shock and be fatal if untreated [53]. This dataset includes more than 120,000 frontal chest X-rays, each of which has been precisely annotated. Employing the data partitioning method described by [19], we conducted tests on the test set.
@@
NIH ChestXray14 [54]. This dataset contains 112,120 chest X-ray images, covering 14 different potentially concurrent chest diseases. For the classification task, we use the official data split for evaluation. For the localization task, we evaluate all 8 categories with localization annotations and follow the settings of [55] to ensure a fair comparison.
@@
ChestX-Det10 [56]. This dataset is a subset of the NIH ChestXray14 dataset, consisting of 3,543 chest X-ray images. These images have been annotated with bounding boxes for 10 diseases/abnormalities by three board-certified radiologists. In this study, we followed the official data splits and conducted qualitative visual evaluations of our model on the test set.
@@
Retinal Fundus. This dataset was compiled from a total of 26,028 retinal fundus images and paired reports from 11,372 individuals. All fundus images were captured using a Zeiss fundus camera and annotated by retinal experts with 10 years of experience for nine common retinal diseases. We divided the data set into training, validation, and testing sets by 8:1:1 based on the patient’s unique ID, which ensures that the same patient does not appear in the same data set. Our model was trained and validated on the training and validation sets, respectively, and classification performance evaluations were done on the testing dataset. To evaluate the performance of our model in pathology localization, we utilized an additional dataset for validation, which was annotated at the pixel level by retinal specialists. Specifically, this dataset includes 437 images of Choroidal neovascularization, 160 images of drusen, and 498 images of retinal hemorrhage.

### v7

@@ Results
This section aims to enhance AFLoc’s clinical applicability by presenting results obtained through leveraging limited labeled data for model fine-tuning on localization tasks.
As expected, AFLoc demonstrates enhanced segmentation performance with increased fine-tuning using labeled data on the SIIM dataset (Supplementary Table9). The Dice score is increased from 0.772 with 1% of the data for fine-tuning to 0.809 with 100% of the data. Furthermore, AFLoc consistently achieves higher Dice scores than all comparative methods across varying portions of labeled fine-tuning data. These results indicate that AFLoc can effectively use limited annotations to achieve enhanced segmentation performance.
The results for pathology localization on the NIH ChestXray14 dataset are depicted in Supplementary Table10. Consistent performance enhancement is achieved by AFLoc over the two comparative methods across different IoU thresholds. At an IoU of 0.3, AFLoc attains a mean localization accuracy of 0.62, surpassing the performance of the two comparative methods (Li et al. (0.49) and Wang et al. (0.22)). Notably, the model excels even under stricter thresholds like IoU = 0.5 and IoU = 0.7, especially for more challenging pathologies such as atelectasis and infiltration.
Extended Data Fig. 7 presents visual examples of AFLoc’s localization performance across four datasets: JSRT, SIIM, ChestX-Det10, and Retinal Fundus. These examples cover chest X-rays with common tiny lesions and retinal hemorrhage images. The predicted regions show strong alignment with the ground-truth annotations, demonstrating AFLoc’s ability to enhance localization accuracy using only limited labeled data.
@@
In this section, we explore the impact of incorporating text features at different levels of granularity – word-level, sentence-level, and report-level – on the performance of AFLoc. The results are listed in Supplementary Table11. When using only one level of text features, the performance of AFLoc deteriorates a lot, especially for word-level and report-level features. One possible explanation could be that the word-level features lack sufficient contextual information to capture the characteristics of the pathology, whereas the report-level features may fail to capture critical diagnostic details due to their overly macroscopic perspective. On the other hand, AFLoc using sentence-level features achieves relatively higher performance, suggesting that sentence-level features can provide richer contextual information, which is important for pathology localization in medical imaging. When these different levels of text features are combined in use, clear performance enhancement is observed, and the best performance is achieved when all three levels are employed, fully supporting our hypothesis that the combination of multi-level information is crucial for improving the accuracy of pathology localization in medical images.
@@
Text encoders play a crucial role in deeply understanding medical concepts and achieving accurate semantic alignment. To investigate the impact of text encoding on pathology localization performance, we implemented the text encoder from the latest large language and vision assistant for biomedicine (LLaVA-Med [li2023llavamed]). LLaVA-Med has shown improvements in open biomedical question-answering tasks. However, as demonstrated in Supplementary Table12, its performance falls short compared to BioClinicalBERT [alsentzer2019publicly]. This discrepancy can be attributed to the superior performance of language models trained on domain-specific datasets (e.g., biomedical and clinical data) tailored for their respective tasks [peng2019transfer]. LLaVA-Med was trained on the PMC-15M dataset, which is extracted from scientific publications in PubMed Central, whereas BioClinicalBERT was trained on electronic health records of intensive care unit patients, closely aligning with the clinical reports used in this study. This domain-specific alignment likely provides BioClinicalBERT an advantage in comprehending intricate medical terminology and improving its localization performance.
@@ Discussion
Pathology localization and diagnosis are essential components of clinical applications in medical imaging. However, existing methods encounter various challenges, notably stemming from the scarcity of annotated data related to generalization [wu2023medklip, tiu2022expert]. In this study, we introduce AFLoc, an annotation-free deep learning model designed to enhance pathology localization by using both clinical reports and imaging data in the pre-training phase.
@@
AFLoc employs a multi-level semantic alignment strategy to enhance the medical feature representation. Given that pathologies often occupy only small regions within images, accurate localization can be challenging for global models. While some methods align global and local text-image features [huang2021gloria], they may miss critical semantic context by focusing excessively on individual words. In contrast, AFLoc adopts a more comprehensive approach by incorporating features at different levels of granularity (word-level, sentence-level, and report-level) during pre-training. This allows AFLoc to capture the necessary details for precise pathology localization.
@@
We evaluated AFLoc’s localization and diagnostic capabilities in an annotation-free setting across three major clinical imaging modalities: chest X-rays, retinal fundus, and histopathology. In the chest X-ray modality, AFLoc demonstrated improved localization performance compared to existing methods, such as MedKLIP and BioViL, across multiple evaluation metrics, including Intersection over Union (IoU), Dice similarity coefficient, and Contrast-to-Noise Ratio (CNR). AFLoc effectively localized a variety of thoracic pathologies, including pneumonia, atelectasis, and pleural effusion, on datasets such as RSNA Pneumonia, MS-CXR, and CheXlocalize. Notably, AFLoc showed good performance in localizing subtle and low-contrast pathologies, such as airspace opacity and atelectasis, which are often difficult to detect but crucial for early diagnosis and treatment planning. Furthermore, AFLoc surpassed the human benchmark on the CheXlocalize dataset across five key tasks, implying its potential to compete with conventional methods that depend on human-reviewed data. Additionally, AFLoc performed well on the COVID Rural dataset, which provides X-rays of patients affected with COVID-19 pneumonia, highlighting its potential for generalizing to emerging diseases. This capability is particularly important for timely and accurate disease detection, especially during global health crises like the COVID-19 pandemic.
@@
The use of AFLoc with retinal fundus images underscores its versatility. In retinal pathology, AFLoc performed well at localizing and identifying features such as choroidal neovascularization and drusen, surpassing other methods like GLoRIA. Given the subtlety and diffuseness of retinal diseases, precise localization is particularly challenging. AFLoc’s adept performance without human-annotated data highlights its potential to improve ophthalmic diagnostics, especially in settings where large-scale data annotation is not feasible.
@@
Furthermore, we illustrated that AFLoc’s performance could be further improved with a small amount of annotated data, demonstrating its scalability and adaptability for clinical applications. While annotation-free localization is a valuable property, fine-tuning with labeled data can elevate AFLoc’s performance. This flexibility renders AFLoc fitting for diverse clinical environments, spanning well-resourced hospitals with access to large labeled datasets to smaller clinics in resource-limited settings. AFLoc’s adeptness in achieving commendable results with minimal annotated data makes it a potentially practical solution for healthcare systems globally, particularly in regions where access to expert annotators is limited.
Finally, to evaluate the model’s impact on diagnostic workflow, we conducted a controlled experiment comparing diagnostic accuracy and reading time under conditions with and without AI support. The study involved two board-certified radiologists, both with more than seven years of diagnostic experience. Data were sourced from the radiologist-annotated test subset of the MIMIC-CXR dataset. A total of 100 chest X-ray images were randomly selected and evenly assigned to two groups: one interpreted by unaided radiologists (Radiologist), and the other by AI-assisted radiologists (Radiologist + AI). In the AI-assisted group, the radiologists were provided with additional information indicating the predicted disease categories and their corresponding localization. As illustrated in Fig. 4, within the clinical workflow, these maps were displayed side by side with the original chest X-ray images, helping radiologists quickly focus on potential abnormalities while retaining full control over diagnostic decisions. Under both conditions, the radiologists were asked to determine the presence or absence of 13 disease categories, along with a “no findings” label, resulting in 14 binary classification tasks per image. One point was awarded for each correct classification, with a maximum score of 14 per case. As shown in Extended Data Fig. 9, the AI-assisted group achieved higher diagnostic scores, with the average score increasing from 10.80 to 11.74. Reading time was also reduced, with the average duration decreasing by 20.5% (from 27.92 seconds to 22.20 seconds). These results indicate that the introduction of AI not only improved diagnostic accuracy but also enhanced reading efficiency, highlighting its potential as a decision-support tool in radiology.
Despite AFLoc’s encouraging performance in zero-shot generalization, several limitations remain and motivate future work. First, the model architecture can be further improved with advanced designs. For instance, a hierarchical multi-scale feature-fusion framework—one that progressively aligns extracted image features with text features at each downsampling scale—could further strengthen feature extraction. In future work, we will further explore these fusion strategies to further elevate AFLoc’s generalization capacity and diagnostic performance.
Second, AFLoc currently generates localization results conditioned on positive classification outputs, which aligns with typical clinical workflows where localization is most relevant for suspected or confirmed findings. In the future, we also must investigate the availability of localization information in cases where the classification is uncertain or incorrect. Specifically, future work will explore integrating uncertainty estimation techniques to quantify the confidence of both classification and localization outputs. This can help flag low-confidence cases for clinician review, even when a heatmap is present. In such scenarios, we plan to include visual or textual alerts indicating potential ambiguity in the prediction. Additionally, future research will consider incorporating clinical prior knowledge into the model to assist in correcting classification errors or inaccurate localizations. We also envision a reinforcement learning framework where clinician feedback on mislocalized or misclassified examples can be used to fine-tune the model, enhancing its robustness over time iteratively.
Third, while our multimodal vision–language model demonstrates promising generalizability across X-ray, retinal fundus images, and histopathology, its adaptability to additional clinical modalities remains to be thoroughly investigated. Modalities such as ultrasound and magnetic resonance imaging—as well as non-visual data types like genomics and electrocardiography—present distinct challenges due to their heterogeneous data characteristics. Incorporating modality-specific encoders may better capture cross-modal semantic alignments. Future work could explore advanced multi-way attention mechanisms or the construction of a unified multimodal representation space, thereby enabling the model to handle three or more diverse modalities, capture complex nonlinear dependencies across heterogeneous data sources.
@@ Methods
Text Encoding In clinical practice, medical texts may originate from various types of content, including clinical observations, diagnostic descriptions, and related information. These texts are typically composed of multiple sentences or phrases. In GLoRIA [huang2021gloria], a block tokenization technique was used, constructing complete words by aggregating multiple subwords. On the other hand, BioViL [boecking2022making] employs a custom dictionary developed from multiple datasets, aiming to reduce the frequency of words being split into subwords. Treating words as independent entities may overlook the complete semantics, leading to incorrect semantic alignment between text and image. In our proposed AFLoc, text features are extracted at three levels: word-level, sentence-level, and report-level, aiming to represent report content more comprehensively and precisely through multi-granularity semantics.
Specifically, we employ BioClinicalBERT [alsentzer2019publicly] as our text encoder to extract text features. When processing a medical report containing words and sentences, each word is tokenized into tokens. The tokenizer inputs the tokenized report into the text encoder , where H is the maximum token length. The text encoder outputs , L is the number of layers, and indicates the feature dimension. We take the average feature of the last 4 layers as the subword-level feature . Then, we aggregate features at the word, sentence, and report levels. Particularly, for a word, all its subword-level features are summarized to obtain word-level features . On the other hand, all corresponding subword-level features are averaged to obtain sentence-level features and one report-level feature .
Image Encoding We adopt the widely used ResNet-50 architecture [he2016deep] as the backbone of our image encoder . Given an input image , we extract image features at different levels from the encoder to represent shallow local features, deep local features, and global features respectively. Specifically, shallow local features are extracted from the third down-sampling stage, while deep local features are extracted from the fourth down-sampling stage. Here, denotes the number of sub-regions in the shallow local features, and represents the feature dimension. The output of the final convolutional layer of the image encoder is average pooled to produce the global image features . To match the dimensionality of the image features with that of the text features, we applied a projection layer to all three levels of image features respectively.
@@
Annotation-Free Pathology Localization and Diagnosis We transform the task into a text-image matching problem using text descriptions, offering greater flexibility as the text descriptions can be easily generated to meet various clinical requirements. By default, we use the automatically generated descriptions based on the pathology’s name and a template for testing. That is, for example, ‘Findings suggesting {lesion}’ for a positive description and ‘No findings suggesting {lesion}’ for a negative description. We refer to such description as ‘Simple description’. In addition, AFLoc also supports the ‘Precise description’ of the pathology, such as ‘Large right-sided pneumothorax, air space opacity in a right infrahilar location’. This ‘precise description’ comes from clinicians with their observation of the images. They can write corresponding descriptions to validate their hypotheses when the diagnosis is uncertain. With the text descriptions, the text features can be obtained with the pre-trained text encoder.
@@
Experimental Details We pre-train AFLoc on MIMIC-CXR, Retinal Fundus, and Quilt-1M dataset for chest X-ray, retinal fundus, and histopathology respectively. All images were uniformly resized to pixels and we followed the same image preprocessing strategy as GLoRIA [huang2021gloria].
A random shuffling and a sentence sampling strategy were applied to enhance the text data. Notably, the reports for the chest X-ray pre-train include multiple sections with different contents. We use the texts within the ’Findings’ and ’Impressions’ sections to train our model. Furthermore, we applied an additional sentence augmentation strategy to enhance the diversity of medical reports for retinal fundus data. Specifically, we generate five semantically equivalent but differently expressed sentences for each sentence by ChatGPT. These equivalent sentences will be randomly replaced during pre-training. For the histopathology dataset, we directly use the corresponding caption provided by the dataset for pre-training.
During pre-training, we used the Adam optimizer with an initial learning rate of 0.00002. A linear learning rate decay mechanism is employed by a decay factor of 0.9. The other hyperparameters were optimized with a corresponding validation dataset.
Evaluation Metrics For the localization task, the Intersection over Union (IoU) and Dice coefficients were used to evaluate the thresholded results. We report the average results at multiple thresholds following the settings from BioViL. For the CheXlocalize dataset, we followed the methodology from [saporta2022benchmarking] that determined an optimal threshold from the validation set and applied it to the test set. Additionally, we used the CNR [boecking2022making] for measuring the difference in heatmap scores between inside and outside the bounding box area, eliminating the need for a hard threshold. This assessment of local similarity is significant, as certain clinical downstream applications might benefit more from heatmap visualizations than discrete segmentation masks. To directly assess localization performance without the influence of classification errors, we reported localization performance on the positive slices of the dataset (those with ground truth segmentations or bounding boxes annotated by radiologists). This approach ensured that the evaluation focused on the accuracy of localization, without being confounded by classification errors [saporta2022benchmarking]. The averages of these metrics are reported on the test set, based on over 1,000 bootstrap repetitions, accompanied by the 95% confidence interval (CI), which was determined using the 2.5th and 97.5th percentiles of the empirical distribution.
@@
We evaluated our model across three different imaging modalities: chest X-ray, retinal fundus, and histopathology. For the chest X-ray experiments, we pre-trained our model on the MIMIC-CXR dataset, then evaluated it across 8 widely used external datasets: RSNA Pneumonia (RSNA), MS-CXR, CheXlocalize, COVID Rural, SIIM-ACR Pneumothorax (SIIM), NIH ChestXray14, ChestX-Det10, and CXR-LT. These datasets encompass 34 types of diseases, such as pneumonia, pleural effusion, and pneumothorax. For the retinal fundus experiments, we pre-trained and evaluated the model using a real-world in-house dataset. For the histopathology experiments, we pre-trained our model on the Quilt-1M dataset and then evaluated it on three external datasets: SICAPv2, WSSS4LUAD, and DHMC LUAD. The details of each dataset are provided below.
@@
MIMIC-CXR [johnson2019mimic]. The MIMIC-CXR dataset is an extensive, publicly available collection of chest X-ray images, paired with free-text radiological reports. It contains 377,110 images that are linked to 227,835 radiological studies. Each study within the database may consist of one or more images, providing a diverse range of scanning perspectives for analysis and research purposes.
@@
RSNA Pneumonia (RSNA) [shih2019augmenting]. The RSNA Pneumonia dataset presents a challenge in identifying visual signals of pneumonia from medical images. Pneumonia typically presents as one or multiple regions of increased opacity in chest X-rays [franquet2018imaging]. Images in this dataset are accompanied by bounding box labels that indicate the areas affected by pneumonia. Over 1,000 cases in this dataset are positive, with each image potentially containing one or more boundary boxes highlighting the pneumonia regions.
@@
MS-CXR [boecking2022making]. MS-CXR is a curated database that features 1,153 image-sentence pairs of bounding boxes along with their corresponding phrases. This collection encompasses 8 different types of cardiopulmonary radiological findings. All the data have been annotated and validated by professional radiologists, ensuring a gold standard for phrase grounding assessment. All data in MS-CXR have been removed from the MIMIC-CXR training set to ensure it was not used in the training process.
@@
CheXlocalize [saporta2022benchmarking]. CheXlocalize is a comprehensively curated chest X-ray dataset that provides high-quality, multi-label, pixel-level segmentations. This dataset includes 234 images, encompassing 643 segmented pathologies. Moreover, CheXlocalize also provides human benchmark segmentations, drawn by a separate group of 3 board-certified radiologists, which can be used to study human benchmark localization performance.
@@
COVID Rural [desai2020chest, tang2020deep]. The COVID Rural project includes 221 chest X-ray images from 105 patients who tested positive for COVID-19. In these images, opaque areas associated with COVID-19 infection are delineated by polygons that connect specific points. The manual annotations of these images were generated by primary physicians and reviewed by experienced radiologists.
@@
SIIM-ACR Pneumothorax (SIIM) [siim]. The SIIM is a widely used pneumothorax dataset designed for the development and evaluation of detection models. In severe cases, pneumothorax can cause obstructive shock and be fatal if untreated [sahn2000spontaneous]. This dataset includes more than 120,000 frontal chest X-rays, each of which has been precisely annotated. Employing the data partitioning method described by [wu2023medklip], we conducted tests on the test set.
@@
NIH ChestXray14 [wang_chestx-ray8_2017]. This dataset contains 112,120 chest X-ray images, covering 14 different potentially concurrent chest diseases. For the classification task, we use the official data split for evaluation. For the localization task, we evaluate all 8 categories with localization annotations and follow the settings of [li2018thoracic] to ensure a fair comparison.
@@
ChestX-Det10 [liu2020chestxdet10]. This dataset is a subset of the NIH ChestXray14 dataset, consisting of 3,543 chest X-ray images. These images have been annotated with bounding boxes for 10 diseases/abnormalities by three board-certified radiologists. In this study, we followed the official data splits and conducted qualitative visual evaluations of our model on the test set.
CXR-LT [CXR-LT, CXR-LT-MIA, PhysioNet]. This dataset is a large-scale chest X-ray dataset designed for studying multi-label and long-tailed disease classification from the CXR-LT 2024 Challenge. Compared to CXR-LT 2023 [holste2023cxr], it includes 19 additional rare disease labels. In this study, we evaluated the zero-shot classification performance of our model on the test sets, specifically for Task 2 (26 labels, including 12 newly annotated ones) and Task 3 (5 unseen labels).
@@
Retinal Fundus. This dataset was compiled from a total of 26,028 retinal fundus images and paired reports from 11,372 individuals. All fundus images were captured using a Zeiss fundus camera and annotated by retinal experts with 10 years of experience for nine common retinal diseases. We divided the data set into training, validation, and testing sets by 8:1:1 based on the patient’s unique ID, which ensures that the same patient does not appear in the same data set. Our model was trained and validated on the training and validation sets, respectively, and classification performance evaluations were done on the testing dataset. To evaluate the performance of our model in pathology localization, we used an additional dataset for validation, which was annotated at the pixel level by retinal specialists. Specifically, this dataset includes 437 images of choroidal neovascularization, 160 images of drusen, and 498 images of intraretinal hemorrhage. The study was approved by the institutional ethics committee. It was in compliance with the principles of the Declaration of Helsinki, the Good Clinical Practice (GCP), the ICH-GCP, and other relevant national laws and regulations.